The difference between two numbers is 19. Half the larger number plus twice the smaller number is 12. What are the numbers?

This is the equation I came up with but I can't make it work.
x-y=19
1/2x+2y=12

Answer by Alan3354(69443)

The difference between two numbers is 19. Half the larger number plus twice the smaller number is 12. What are the numbers?

This is the equation I came up with but I can't make it work.
x-y=19
1/2x+2y=12 x2 --> x + 4y = 24
----------
x - y = 19
x +4y = 24
------------ Subtract
-5y = -5
y = 1
x = 20